OpenStudy (anonymous):

the difference between five and the product of seven and a number equal to nine.

Parth (parthkohli):

Let's say that the number is \(x\):\[7x - 5 = 9 \]Solve this equation. Hint: Start by adding 5 to both sides.

OpenStudy (ghazi):

i guess it's 5-7x=9, so x would be you can calulate

Parth (parthkohli):

No, @ghazi You may apply logic and see how it can't be that.

OpenStudy (ghazi):

because you can't have 9 by my equation

Parth (parthkohli):

No, apply your conscience and look how you're getting such a complicated answer for your equation:P

OpenStudy (ghazi):

if you'll substitute x= -4/9 you'll have 9 @ParthKohli

OpenStudy (ghazi):

well i just went on the language and its sequence

Parth (parthkohli):

Actually you can have two answers here.

OpenStudy (ghazi):

and you tried to have the easier one so you just wrote equation in a straight forward manner

Parth (parthkohli):

Ghazi, I know what you mean, but you must see how it's a little tedious to get such a solution.

Parth (parthkohli):

7x should be greater than 5 according to the intuition, so you'd make it \(7x - 5 = 9\).

OpenStudy (ghazi):

okay :) let her accept what she thinks is easier one....

OpenStudy (ghazi):

well if you'll substitute x=-4/7 which is also greater than 7*x as per your solution still it will be in accordance

Parth (parthkohli):

Well, the intuition says so.